Institut: MATEMATIK OG DATALOGI

Kursustitel: CRYPTOLOGY

Semesteromfang/timetal: 0.2 semester / 2 hours lecture, plus 2 hours discussion section per week.

Indgangskrav: DM02 and either DM11 or MM05.

Målsætning: The purpose of this course is to study cryptology, which is cryptolgraphy plus cryptanalysis - the creation of secret codes and the possiblities for breaking them. We will also study cryptographic protocols for the security of information. Many of the newer cryptosystems and cryptographic protocols are based on number theoretic problems, so these number-theoretic problems and algorithms for them will also be discussed in this course.

Cryptology has many applications including sending private messages, enabling commerce over the Internet (through encryption of credit card numbers, electronic money, secure methods for electronic signatures on documents, etc.), authentication such as PIN codes for Dankort and logins, and secret sharing (requiring that k out of m people participate before some sensitive action can occur).

Emneoversigt: Public-key cryptography, RSA, signature schemes, authentication, secret sharing, zero-knowledge proofs, primality testing, factoring, discrete logarithms.

Evaluering: 4 hour written exam with external examiner. The grading will be on the 13 point scale. In addition, there will be a few homework assignments, which must be approved in order to take the exam.

Undervisningsform: Lectures by the teacher.

Lærer: Joan Boyar.

Litteratur: Douglas R. Stinson, Cryptography. Theory and Practice, CRC Press, 1995.